Two people are 3 1/2 miles apart walking towards each other. Their uniform walking differ by 1 mph. They will meet in 1/2 hour. What is the uniform rate of the slower walker? how much of the distance will he/she travel?
Answer by stanbon(75887)

Two people are 3 1/2 miles apart walking towards each other. Their uniform walking differ by 1 mph. They will meet in 1/2 hour. What is the uniform rate of the slower walker? how much of the distance will he/she travel?
----
Faster walker DATA:
rate = r mph ; time = 1/2 hour ; distance = (1/2)r miles
-----
Slower walker DATA:
rate = r-1 mph ; time = 1/2 hour ; distance = (1/2)(r-1) miles
----
Equation:
(1/2)r + (1/2)(r-1) = 7/2
---
r + r -1 = 7
2r = 8
r = 4 mph
r-1 = 3 mph (rate of the slower walker.
distance for slower walker: (1/2)3 = 3/2 miles
===================================
Cheers,
Stan H.
===================================